OpenStudy (anonymous):

find the difference in 4ab2 − 3ab2

OpenStudy (anonymous):

\[4ab^2 - 3ab^2 = (4 - 3)ab^2 = ab^2\]

OpenStudy (anonymous):

this equals ab^2. just subtract when there are like terms